Operations Report – February 26th

Crew 222 Operations Report 02/26/2020

SOL: 10

Name of person filing report: Valentin PRUDHOMME

Non-nominal systems: N/A

Notes on non-nominal systems: N/A

Tuesday generator:
Off at 5:30am
On at 8:45pm SOC was 80%

Wednesday generator:
Off at 6:00am
On at 9:00pm SOC was XX%

Diesel Reading – 72%

Station Propane Reading – 84% (delivered this morning)

Ethanol Free Gasoline – 0.5 gallons

Water (loft tank): 47 gallons

Water Meter: 149505.59 units

Water (static tank) – 258 gallons

Static to Loft Pump used – yes

Water in GreenHab – 47.5 gallons

Water in ScienceDome: 0 gallons

Toilet tank emptied: no

Deimos rover used: yes (also moved by Dr Rupert)
Hours: 146.3
Beginning charge: 100
Ending charge: 97%
Currently charging: yes

Sojourner rover used: ASSIGNED TO DIRECTOR
Hours: N/A
Beginning charge: N/A
Ending charge: N/A
Currently charging: N/A

Spirit rover: no
Hours: 132.8
Beginning charge: N/A
Ending charge: N/A
Currently charging: yes

Opportunity rover used: moved from campus to the shop to repair the brakes. Hours: N/A
Beginning charge: N/A
Ending charge: N/A
Currently charging: yes

Curiosity rover used: yes
Hours: 138.8
Beginning charge: 100%
Ending charge: 94%
Currently charging: yes

Notes on rovers: N/A

ATV’s Used: yes
Reason for use: Run during the engineer EVA for the battery. Oil Added? no
ATV Fuel Used: N/A
Hours the ATVs were Used today: 10 minutes
Notes on ATVs: The red one was jumped bu Dr Rupert during the afternoon. One of the blue ones doesn’t start. We will try to charge it during the night thanks to the battery charger of the RAM.

HabCar used and why, where? The habcar is now in the shop awaiting repairs.

CrewCar used and why, where? N/A

General notes and comments: N/A

Summary of internet: All ok.

Summary of suits and radios: Suits 7, 8, 9, 10 were used during the EVA. The suit number 7 was working well at the beginning but one hour after the beginning of the EVA the fan was stopping at random moments but was restarting after. The fan never stopped totally but wasn’t working optimally.

Summary of Hab operations: N/A

Summary of GreenHab operations: N/A

Summary of ScienceDome operations: N/A

Summary of RAM operations: N/A

Summary of any observatory issues: N/A

Summary of health and safety issues: N/A

Questions, concerns and requests to Mission Support: N/A

Copyright © The Mars Society. All rights reserved. | Main Site